# TopoTune: a framework for generalized combinatorial complex neural networks

### v2

@@ Appendix B Software
Remark. We decided to design the software module of TopoTune, i.e., how to implement GCCNs, as we did for mainly two reasons: (i) the full compatibility with TopoBenchmark (implying consistency of the combinatorial complex instantiations and the benchmarking pipeline), and (ii) the possibility of using GNNs as neighborhood message functions that are not necessarily implemented with a specific library. However, if the practitioner is interested in entirely wrapping the GCCN implementation into Pytorch Geometric or DGL, they can do it by noticing that a GCCN is equivalent to a heterogeneous GNN where the heterogeneous graph the whole augmented Hasse graph, with node types given by the rank of the cell (e.g. 0-cells, 1-cells, and 2-cells) while the edge type is given by the per-rank neighborhood function (e.g. "0-cells to 1-cells" or "2-cells to 1-cells" for $\mathcal{N}^{0}_{I,\uparrow}$ and $\mathcal{N}^{2}_{I,\downarrow}$, respectively).
Algorithm 1 TopoTune
1:Class TopoTune(torch.nn.Module):
2:procedure Init(neighborhoods, $\omega_{n}$, $\omega_{n}\_params$, layers)
3:     $self.omega\_n\_submodels\leftarrow[]$
4:     for $l\leftarrow 1$ to layers do
5:         $layer\_models\leftarrow[]$
6:         for each $nb$ in neighborhoods do
7:              $model\leftarrow\omega_{n}(\omega_{n}\_params)$
8:              $layer\_models.append(model)$
9:         end for
10:         $self.omega\_n\_submodels.append(layer\_models)$
11:     end for
12:end procedure
13:procedure Forward($x$)
14:     for each layer in $self.omega\_n\_submodels$ do
15:         $outputs\leftarrow[]$
16:         for each $\omega_{n}\_model$ in layer do
17:              $hasse\_graph\leftarrow self.expand\_to\_strictly\_aug\_hasse\_graph(x)$
18:              $outputs.append(\omega_{n}\_model(hasse\_graph))$
19:         end for
20:         $x\leftarrow self.aggregate\_rank\_wise(outputs)$
21:     end for
22:     return $x$
23:end procedure
24:
25:Example Instantiation:
26:$neighborhoods\leftarrow[[[0,0],up\_adjacency],[[2,1],incidence]]$
[…]
28:$\omega_{n}\_params\leftarrow\{num\_layers:2,heads:4\}$
29:$layers\leftarrow 4$
30:$model\leftarrow\text{TopoTune}(neighborhoods,\omega_{n},\omega_{n}\_params,layers)$
Appendix C Additional details on experiments
@@ C.2 Datasets
C.3 Hyperparameter search
@@ C.4 Hardware
Appendix D Model Size
@@
Appendix E Performance versus Model Size
We show the plots similar to Fig. 5 for all datasets. Again here, the best model determines the amount of GCCN layers and GNN sublayers we keep constant.
Figure 6: Performance versus Size, scaled to best-performing model.

### v3

@@ Appendix D Software
Remark. We decided to design the software module of TopoTune, i.e., how to implement GCCNs, as we did for mainly two reasons: (i) the full compatibility with TopoBenchmark (implying consistency of the combinatorial complex instantiations and the benchmarking pipeline), and (ii) the possibility of using GNNs as neighborhood message functions that are not necessarily implemented with a specific library. However, if the practitioner is interested in entirely wrapping the GCCN implementation into Pytorch Geometric or DGL, they can do it by noticing that a GCCN is equivalent to a heterogeneous GNN where the heterogeneous graph the whole augmented Hasse graph, with node types given by the rank of the cell (e.g. 0-cells, 1-cells, and 2-cells) while the edge type is given by the per-rank neighborhood function (e.g. ”0-cells to 1-cells” or ”2-cells to 1-cells” for $\mathcal{N}^{0}_{I,\uparrow}$ and $\mathcal{N}^{2}_{I,\downarrow}$, respectively).
Appendix E Additional details on experiments
@@ E.2 Datasets
Dataset selection and limitations
The datasets employed in this work and other TDL studies are predominantly adapted from the GNN literature. Among these, molecular datasets stand out due to the inherent importance of cycles and hyperedges, which effectively capture chemical rings and functional groups. These are structures that are naturally represented in topological domains.
While TDL methods are not intrinsically constrained to these datasets, the lifting procedures used to construct higher-order cells introduce computational bottlenecks, particularly in memory usage. For instance, operations such as cycle detection and clique enumeration, required for constructing cellular complexes or simplicial complexes, respectively, become computationally prohibitive for large or densely connected graphs.
To address these limitations, ongoing research is focused on developing scalable lifting procedures that can extend TDL methods to broader datasets, including those with more complex structures or larger scales. For example, Bernárdez et al. (2024) propose innovative topological liftings, paving the way for more scalable and applicable datasets in TDL.
E.3 Hyperparameter search
@@ E.4 Hardware
Appendix F Iterating upon existing CCNNs
We design and test GCCNs that replicate the neighborhood structure and training scheme of existing CCNNs, and replace their message-passing schemes with the same GNNs considered in Table 1. Results in Table 3 show that this sometimes leads to better performance on benchmark datasets.
Model MUTAG PROTEINS NCI1 NCI109 Cora Citeseer PubMed
SCCN (Yang et al., 2022)
Benchmark results (Telyatnikov et al., 2024) 70.64 ± 5.90 74.19 ± 2.86 76.60 ± 1.75 77.12 ± 1.07 82.19 ± 1.07 69.60 ± 1.83 88.18 ± 0.32
CWN (Bodnar et al., 2021a)
Benchmark results (Telyatnikov et al., 2024) 80.43 ± 1.78 76.13 ± 2.70 73.93 ± 1.87 73.80 ± 2.06 86.32 ± 1.38 75.20 ± 1.82 88.64 ± 0.36
Appendix G Model Size
@@
Appendix H Model Training Time
We provide training times for all experiments reported on in Section 6. We measure these training times by running each experiment on a single A30 NVIDIA GPU. We note that these times include the on-the-fly graph expansion method, which slows down the model forward proportionally to dataset size. We plan on moving this process into data preprocessing in the future.
Table 6: Model training time (seconds) corresponding to results reported in Table 1.
Graph-Level Tasks Node-Level Tasks
Cellular
CCNN (Best Model on TopoBenchmark) 100 ± 23 132 ± 19 238 ± 89 254 ± 39 228 ± 44 75 ± 15 57 ± 4.4 128 ± 50
GCCN $\omega_{\mathcal{N}}$ = GAT 80 ± 11 64 ± 10 778 ± 118 486 ± 75 3173 ± 954 46 ± 3 63 ± 1 202 ± 22
GCCN $\omega_{\mathcal{N}}$ = GCN 43 ± 7 67 ± 16 544 ± 40 495 ± 108 4013 ± 620 46 ± 4 65 ± 3 149 ± 12
GCCN $\omega_{\mathcal{N}}$ = GIN 61 ± 18 59 ± 18 523 ± 119 386 ± 76 3301 ± 440 64 ± 8 77 ± 2 207 ± 33
GCCN $\omega_{\mathcal{N}}$ = GraphSAGE 43 ± 12 43 ± 3 691 ± 80 364 ± 102 2863 ± 262 49 ± 2 60 ± 3 211 ± 25
GCCN $\omega_{\mathcal{N}}$ = Transformer 50 ± 19 786 ± 147 1005 ± 27 1484 ± 181 15320 ± 5386 121 ± 20 94 ± 20 5459 ± 1374
GCCN $\omega_{\mathcal{N}}$ = Best GNN, 1 Aug. Hasse graph 33 ± 7 70 ± 24 451 ± 123 441 ± 130 3162 ± 340 47 ± 5 72 ± 6 194 ± 35
Simplicial
CCNN (Best Model on TopoBenchmark) 123 ± 57 104 ± 28 172 ± 50 183 ± 62 178 ± 86 143 ± 16 75 ± 23 114 ± 18
GCCN $\omega_{\mathcal{N}}$ = GAT 25 ± 5 70 ± 17 755 ± 158 794 ± 151 2242 ± 275 49 ± 3 68 ± 2 192 ± 38
GCCN $\omega_{\mathcal{N}}$ = GCN 40 ± 7 138 ± 26 548 ± 185 603 ± 181 2428 ± 833 49 ± 5 67 ± 2 167 ± 22
GCCN $\omega_{\mathcal{N}}$ = GIN 61 ± 7 66 ± 21 904 ± 180 538 ± 39 3603 ± 475 71 ± 6 77 ± 8 210 ± 42
GCCN $\omega_{\mathcal{N}}$ = GraphSAGE 31 ± 3 61 ± 27 572 ± 124 511 ± 74 1721 ± 201 51 ± 3 74 ± 8 221 ± 37
GCCN $\omega_{\mathcal{N}}$ = Transformer 35 ± 5 947 ± 333 1386 ± 404 1360 ± 410 7979 ± 1373 146 ± 58 77 ± 2 5281 ± 827
GCCN $\omega_{\mathcal{N}}$ = Best GNN, 1 Aug. Hasse graph 25 ± 2 78 ± 27 598 ± 31 312 ± 7 2681 ± 910 52 ± 4 72 ± 8 156 ± 16
Hypergraph
CCNN (Best Model on TopoBenchmark) 127 ± 48 96 ± 20 220 ± 74 128 ± 49 387 ± 105 121 ± 38 48 ± 1 177 ± 71
Table 7: Model training times (seconds) corresponding to results in Table 3.
Model MUTAG PROTEINS NCI1 NCI109 Cora Citeseer PubMed
SCCN (Yang et al., 2022)
Benchmark results (Telyatnikov et al., 2024) 11 ± 2 60 ± 18 247 ± 65 311 ± 83 102 ± 39 101 ± 41 143 ± 35
GCCN, on ensemble of strictly aug. Hasse graphs *2, dig 14 ± 1 75 ± 8 413 ± 120 298 ± 15 121 ± 2 172 ± 6 285 ± 20
GCCN, on 1 aug. Hasse graph *2, dig 5 ± 1 59 ± 10 283 ± 90 217 ± 100 110 ± 3 166 ± 10 376 ± 27
CWN (Bodnar et al., 2021a)
Benchmark results (Telyatnikov et al., 2024) 11 ± 2 43 ± 5 240 ± 50 252 ± 92 54 ± 25 52 ± 5 119 ± 14
GCCN, on ensemble of strictly aug. Hasse graphs *2, dig 12 ± 1 73 ± 10 536 ± 38 426 ± 90 91 ± 17 49 ± 1 125 ± 19
GCCN, on 1 aug. Hasse graph *2, dig 11 ± 1 62 ± 11 573 ± 107 410 ± 64 96 ± 2 46 ± 1 130 ± 20
Appendix I Performance versus Size Complexity
In, Fig. 10, we extend Fig. 5 to all benchmark datasets. As before, we keep GCCN layers and GNN sublayers in each subplot constant, matching those of the best model of that dataset.
Figure 10: Performance versus size, scaled to best model. We consider models within 10% of the best performance on each dataset.